Total cost =400+20q-2qsquare+2/3qcube, determine the firm's Short run supply curve with its graph

1
Expert's answer
2022-01-13T09:01:44-0500

Solution:

TC = 400 + 20q – 2q2 + 2/3q3


SMC = "\frac{\partial TC} {\partial Q}" = 20 – 4q + 2q2


SAC = "\frac{TC}{Q}" = 400 + 20q – 2q2 + "\frac{2}{3}" q3 "\div" q = "\frac{400}{q}" + 20 – 2q + "\frac{2}{3}"q2


SAVC = "\frac{VC}{Q}" = 20q – 2q2 + "\frac{2}{3}"q3"\div" q = 20 – 2q + "\frac{2}{3}"q2